OpenStudy (unklerhaukus):

\[\lim_{R\to+\infty}\frac{x^2}{\sqrt{R^2+x^2}-R}%&=\\%&=\]

OpenStudy (ikram002p):

multiply with conjucate ?

OpenStudy (ikram002p):

you would got limit \( (\sqrt( R^2+x^2) +R \) which is \(\infty \)

OpenStudy (unklerhaukus):

\[ \lim_{R\to+\infty}\frac{x^2}{\sqrt{R^2+x^2}-R}\\ =\lim_{R\to+\infty}\frac{x^2}{\sqrt{R^2+x^2}-R}\times\frac{\sqrt{R^2+x^2}+R}{\sqrt{R^2+x^2}+R}\\ =\lim_{R\to+\infty}\frac{x^2(\sqrt{R^2+x^2}+R)}{(R^2+x^2)-R^2}\\ = \]

OpenStudy (unklerhaukus):

\[=\lim_{R\to+\infty}\frac{x^2(\sqrt{R^2+x^2}+R)}{x^2}\\ =\lim_{R\to+\infty}\sqrt{R^2+x^2}+R \]

OpenStudy (kirbykirby):

which should just yield \(\infty\)

OpenStudy (unklerhaukus):

hmm, it was meant to become zero,

OpenStudy (unklerhaukus):

whops, i meant \[\begin{align} &= \lim_{R\to+\infty}\frac{x^2}{\sqrt{R^2+x^2}+R} \\ &= \lim_{R\to+\infty}\frac{x^2}{\sqrt{R^2+x^2}+R} \times\frac{\sqrt{R^2+x^2}-R}{\sqrt{R^2+x^2}-R} \\ &=\lim_{R\to+\infty}\frac{x^2(\sqrt{R^2+x^2}-R)}{(R^2+x^2)-R^2} \\ &=\lim_{R\to+\infty}\frac{x^2(\sqrt{R^2+x^2}-R)}{x^2} \\ &=\lim_{R\to+\infty}\sqrt{R^2+x^2}-R %&\leadsto\infty \\ %& %&\neq 0 \end{align}\]

OpenStudy (unklerhaukus):

How can i show this one equals zero

OpenStudy (xapproachesinfinity):

try factoring R^2?

OpenStudy (xapproachesinfinity):

then factor R

OpenStudy (xapproachesinfinity):

you will get -1x0

OpenStudy (kirbykirby):

Well when you got \[ \lim_{R\to+\infty}\frac{x^2}{\sqrt{R^2+x^2}+R}\] You can plug in infnity directly in the denominatory and you get \( \infty + \infty\) which is okay and defined to be \(\infty\), so you essentially have \[ \lim_{R \rightarrow \infty}\frac{x^2}{R}=0\], it behaves like 1/x as x goes to infinity

OpenStudy (kirbykirby):

wait this looks like your original problem but but the denominator changes signs.. hm I don't think I read your full solution correctly xD

OpenStudy (xapproachesinfinity):

yeah actually from the start it is zero no further steps needed

OpenStudy (unklerhaukus):

hmm ok , thanks

OpenStudy (kirbykirby):

well from the very start you have \(\infty - \infty\) in the denominator which is undefined

OpenStudy (xapproachesinfinity):

no he said that is + not negative so it like you wrote x^2/R which tends to 0

OpenStudy (unklerhaukus):

i see now that sign in the denominator changes everything,

OpenStudy (xapproachesinfinity):

yeah any sign would change the things lol

OpenStudy (kirbykirby):

or so it iis + in the denominator, then that is great! :)

OpenStudy (unklerhaukus):

@ikram002p, @kirbykirby , @xapproachesinfinity thanks y'all

OpenStudy (kirbykirby):

:)

OpenStudy (xapproachesinfinity):

yw!

OpenStudy (unklerhaukus):

(i was just trying to understand physics lecture notes, and i had copied a sign wrong, but you've all cleared it up now)